OpenStudy (rebeccapink6398):

OMG! Having trouble... could someone plz help? (I don't want answer) Find the simple interest. principal = $10,400; rate = 12.5%; time = 11 years

OpenStudy (rebeccapink6398):

Could someone please explain it?

OpenStudy (anonymous):

\[\huge SI = \frac {PNR}{100} =\frac {10,400 \times 12.5 \times 11}{100}\] \[\huge SI = \frac {PNR}{100} = {10,4 \times 12.5 \times 11}=14300\] Hence Simple interest = 14300 dollars..

OpenStudy (rebeccapink6398):

@dpasingh ?? plz explain it. I don't get what u put ;(

OpenStudy (rebeccapink6398):

@ganeshie8 PLZ PLZ PLZ PLZ PLZ PLZ PLZ HELP!!! :D

OpenStudy (anonymous):

First change 12.5% into a decimal, 0.125. Then multiply by 10,400 by 0.125 when it says simple interest.

OpenStudy (anonymous):

When it says the total amount saved/owed do the same but also multiply by the number of years and add them together.

OpenStudy (anonymous):

Here P = 10,400 (Principal amount) N= 11 years (time ) R= 12.5 % ( rate of interest)

OpenStudy (anonymous):

@RebeccaPink6398

OpenStudy (rebeccapink6398):

grrr..I still don't get it :(

OpenStudy (anonymous):

12.5% as a decimal is 0.125, so multiply 0.125 x 10,400.

OpenStudy (rebeccapink6398):

=1300

OpenStudy (anonymous):

Good. :) So is that one of the answers?

OpenStudy (rebeccapink6398):

YES :D

OpenStudy (rebeccapink6398):

@Denisekay ... how about this one? How do I do it? Based on simple interest, find the total amount saved or owed. principal = $1056; rate = 3.75%; time = 6 years

OpenStudy (anonymous):

Yay! :D Just make sure if it says total amount multiply the principal by the rate and the time then add them together.

OpenStudy (anonymous):

3.75% = 0.0375

OpenStudy (rebeccapink6398):

@Denisekay I don't get it :(

OpenStudy (anonymous):

Don't get what? :(

OpenStudy (rebeccapink6398):

I don't get how to do it D:

OpenStudy (anonymous):

1,056 x 0.0375 = 39.6 1,056 x 6 = 6336 39.6 + 6336 = 6375.6